# DR-RAG: Applying Dynamic Document Relevance to Retrieval-Augmented Generation for Question-Answering

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has significantly demonstrated the performance of Large Language Models (LLMs) in the knowledge-intensive tasks, such as Question-Answering (QA). RAG expands the query context by incorporating external knowledge bases to enhance the response accuracy. However, it would be inefficient to access LLMs multiple times for each query and unreliable to retrieve all the relevant documents by a single query. We find that even though there is low relevance between some critical documents and query, it is possible to retrieve the remaining documents by combining parts of the documents with the query. To mine the relevance, a two-stage retrieval framework called Dynamic-Relevant Retrieval-Augmented Generation (DR-RAG) is proposed to improve document retrieval recall and the accuracy of answers while maintaining efficiency. Also, a small classifier is applied to two different selection strategies to determine the contribution of the retrieved documents to answering the query and retrieve the relatively relevant documents. Meanwhile, DR-RAG call the LLMs only once, which significantly improves the efficiency of the experiment. The experimental results on multi-hop QA datasets show that DR-RAG can significantly improve the accuracy of the answers and achieve new progress in QA systems.
DR-RAG: Applying Dynamic Document Relevance to Retrieval-Augmented Generation for Question-Answering
Zijian Hei∗,1,2   Weiling Liu∗,1,3   Wenjie Ou∗,1,4   Juyi Qiao1   Junming Jiao1   Zhiqing Zhu1   Guowen Song†,1 1Li Auto Inc.   2Sun Yat-sen University   3Northeastern University, China   4Sichuan University thanks:  Equal contribution to this work during internshiping at Li Auto.    † Corresponding author 1 {songguowen, zhuzhiqing}@lixiang.com
@@ 1 Introduction
Large language models (LLMs) have recently made significant improvement in the field of Natural Language Processing (NLP), especially in text generation tasks Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023b); Anil et al. (2023); Ouyang et al. (2022); Touvron et al. (2023a). Although LLMs perform well in various application scenarios, there are still challenges to the accuracy and timeliness of the generated text, such as in real-time domains. LLMs with intrinsic parameter memories may generate inaccurate or even incorrect text when faced with up-to-date query Min et al. (2023); Mallen et al. (2022); Muhlgay et al. (2023). The problem is known as hallucination, where the text generated by LLMs fail to correspond to real-world knowledge Ji et al. (2023); Zhang et al. (2023); Kwiatkowski et al. (2019). Therefore, Retrieval-Augmented Generation (RAG) frameworks are proposed to improve the accuracy of generated text by combining relevant information from external knowledge base with query Arora et al. (2023); Lewis et al. (2020); Borgeaud et al. (2022). RAG has effectively demonstrated its superiority in knowledge-intensive tasks such as open-domain Question-Answering (QA) and has achieved new progress in the LLMs’ performance.
Figure 1: QA performance (F1) and time for different RAG frameworks. We use the GPT-3.5-Turbo-Instruct as the base LLM on the multi-hop QA datasets (MuSiQue, HotpotQA and 2Wiki).
@@ 3 Method
Figure 3: An overview of DR-RAG. In step 1, we retrieve static-relevant documents (SR-Documents) due to high relevance with the query. Then we concatenate SR-Documents with the query to retrieve multiple dynamic-relevant documents (DR-Documents) in step 2. Finally, we select each of DR-Documents in turn to splice with the query and SR-Documents and feed them into the classifier to select the most relevant DR-Document.
@@ 3.2 DR-RAG
In this section, we will give a comprehensive description about the DR-RAG framework, a new two-stage retrieval method compared to traditional reranking methods NetEase Youdao (2023); Chen et al. (2024). From Fig. 3, we retrieve $k_{1}$ documents through SM method (first-retrieval stage) and employ a classifier $C$ to model the dynamic relevance between documents (selection process) to enhance the recall rate of the remaining $k_{2}$ documents. The classifier $C$ lies in assessing the dynamic relevance between documents to determine whether the information from the documents is crucial to answer the present query. DR-RAG aims to retrieve $\boldsymbol{d_{stat}^{*}}$ from the $k_{1}$ documents in the first-retrieval stage and then retrieve $\boldsymbol{d_{dyn}^{*}}$ from the $k_{2}$ documents (second-retrieval stage). By combining these key documents, LLMs have the ability to generate the answer of queries more accurately.
@@ 3.2.1 Query Documents Concatenation
As mentioned before, due to the low relevancy between dynamic-relevant documents and the query, the documents are difficult to be retrieved. Moreover, the only relevant information ‘spouse/wife’ between them is also obscured by the mixed information in the knowledge base because too many documents in $D$ will contain ‘spouse’. Therefore, Query Documents Concatenation (QDC) method aims to employ the sentence to match for more useful and relevant information. After the first-retrieval stage, we will obtain $k_{1}$ static-relevant documents and concatenate $\boldsymbol{q}$ with each document to form multiple $<\boldsymbol{q},\boldsymbol{d_{i}}>,i\in k_{1}$ pairs. Moreover, dynamic-relevant documents from $D$ can be retrieved by corresponding $<\boldsymbol{q},\boldsymbol{d_{i}}>,i\in k_{1}$ pair in the second-retrieval stage. As the case in Fig 3, when $\boldsymbol{q}$ and $\boldsymbol{d_{stat}^{*}}$ are concatenated, the query contains both the name ‘Johan Ludvig Heiberg’ and the relationship ‘spouse/wife’, which is essentially similar to $\boldsymbol{d_{dyn}^{*}}$. Therefore, $\boldsymbol{d_{dyn}^{*}}$ is more clearly related to the query and thus easily retrieved. The whole process is:
$\begin{split}\boldsymbol{d_{m}}&=\texttt{Retriever}(\boldsymbol{q}),m\in[1,k_{1}]\\ \boldsymbol{q}^{*}_{m}&=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d_{m}})\\ \boldsymbol{d_{n}}&=\texttt{Retriever}(\boldsymbol{q}^{*}_{m}),n\in[k_{1}+1,k_{1}+k_{2}]\\ answer&=\texttt{LLM}(\texttt{Concat}(\boldsymbol{d_{1}},\boldsymbol{d_{2}},...\boldsymbol{d_{k_{1}+k_{2}}},\boldsymbol{q}))\\ \end{split}$ (1)
where $k_{1}+k_{2}$ is equal to $k$. Retriever is a common SM method. $\boldsymbol{d_{i}},i\in[1,k]$ is the $i$-th relevant document retrieved from $D$. LLM is a large language model to obtain the answer. $answer$ represents the output to answer the query.
@@ 3.2.2 Classifier for Selection
$\begin{split}\texttt{Classifier}(\boldsymbol{q},\boldsymbol{d}^{*},\boldsymbol{d}^{*})&=positive\\ \texttt{Classifier}(\boldsymbol{q},\boldsymbol{d}^{*},\boldsymbol{d}^{\Delta})&=negative\\ \texttt{Classifier}(\boldsymbol{q},\boldsymbol{d}^{\Delta},\boldsymbol{d}^{\Delta})&=negative\\ \end{split}$ (2)
where Classifier represents the binary classification model. $positive$ and $negative$ results indicate whether the two documents are critical for the query.
Classifier Inverse Selection    In this approach, we selectively exclude some irrelevant documents from the retrieved $k$ documents to minimize document redundancy. Specifically, after obtaining $k$ documents in stages, we pair them pairwise as $<\boldsymbol{q},\boldsymbol{d_{m}},\boldsymbol{d_{n}}>$ and get $C_{k}^{2}$ pairs. The pairs, with the current query $\boldsymbol{q}$, are collectively fed into the classifier $C$. Similarly, when the classification result of a document and the remaining $k$-1 documents is negative, then we consider the document as redundant and should be removed. The whole process is:
Table 2: Results on different datasets with Llama3 (8B) as LLM. We emphasize our results in bold.
@@
Classifier Forward Selection    Unlike the CIS method proposed above, CFS method aims to remove the irrelevant dynamic-relevant documents in the second-retrieval stage. To achieve this goal, we search for a document $d_{step2}$ from $D$ according to the $<\boldsymbol{q},d_{step1}>$ pair, and feed both the query and documents into $C$. When the classification result is negative, we will exclude the dynamic-relevant document in the current retrieved documents, and search for the next dynamic-relevant document which can be classified as positive with $d_{step1}$.
@@ 4 Experiment Settings
The experimental details will be described in this section. Due to space constraints, the descriptions of implementation details, retrieval strategy and baseline can seen in Appendix A.2, A.3 and A.4.
@@ 4.1 Dataset
Table 3: Results on different LLMs compared to Adaptive-RAG. We emphasize our best results in bold.
@@
Llama3-8B Query Document Concatenation 25.10 37.10 32.00 46.05 59.98 53.09 45.79 54.62 39.70
@@ 5 Results and Analysis
Table 4: Results of different retrieval strategies. Actual numbers represents the actual document numbers that we feed into LLMs.
@@ 5.1 Main Results
Experimental results effectively demonstrate the effectiveness of our proposed DR-RAG. Table 2 and 3 presents the performance of DR-RAG in answering multihop query, highlighting the advantages of our approach compared to the current RAG framework Jeong et al. (2024); Asai et al. (2024) across multiple metrics. Table 4 shows the performance of DR-RAG across various retrieval strategies.
As shown in Table 2, DR-RAG achieves higher accuracy across multiple datasets compared to other RAG frameworks, which is in line with our expectations. When retrieving the same $k$ documents, DR-RAG can achieve a higher recall rate and a higher percentage of correct answers. The core idea of our work is to provide enough information for LLMs to answer the query. To fully validate the effectiveness of DR-RAG in multiple dimensions, we follow the metrics (F1, EM, Acc) of Adaptive-RAG and choose the same metrics for the evaluation. From the results, DR-RAG achieves better performance than other baseline RAG frameworks (self-RAG and Adaptive-RAG) on all three metrics. Moreover, DR-RAG is also less than other RAG frameworks in terms of the number of LLMs responses and the time consumed in QA systems.
@@ 5.2 Analysis
Effects of Recall Rate    The ability of LLMs to answer domain-specific query correctly almost depends on whether all the necessary information is included in the prompt context. When relevant information is missing, it is difficult for LLMs with the hallucination problem to accurately answer the query. Without relevant information, LLMs may rely on existing fixed knowledge to response the queries and lead to incorrect answers. Therefore, we devise a method to try to improve the retrieval capability and retrieve as enough relevant documents as possible. Table 5 illustrates the answers of the query with and without sufficient information provided to LLMs. As seen in Table 4, in 2Wiki, our retrieval strategy already achieves a recall rate of 98% when top-$k$ is 6. When we feed enough relevant information into LLMs, the accuracy of their answers can be improved accordingly. CFS method achieves higher recall rate by 26.4% and 8.6% than BM25 and SM methods, respectively, which proves the feasibility of DR-RAG.
@@
Increase Recall with Lower Documents    In CFS method, it seems impossible to find a match for every $<\boldsymbol{q},\boldsymbol{d}>$ pair in the second-retrieval stage because the documents we need have been retrieved. Therefore, there will be cases where the total number of our retrieved documents is less than $k$. For instance, in the HotpotQA dataset, when $k$ is set to 6, the average number of documents actually provided to LLMs is only 5.35, thereby reducing irrelevant information to some extent. CFS method in Table 4 achieves a higher recall rate while retrieving fewer actual numbers of documents compared to QDC method. CFS method yields higher scores across the three metrics in our experiments and achieves more significant retrieval capabilities with lower redundant inputs than other methods.
Ablation Study
Table 6: Ablation study on MuSiQue by Llama3-8B.
top-k LLMs EM F1 Acc
3 CFS 48.71 62.42 55.26
w/o QDC 46.26 59.37 52.60
CIS 44.88 57.05 50.94
w/o QDC 44.82 57.00 50.88
4 CFS 48.30 62.81 55.22
w/o QDC 47.78 61.63 54.73
CIS 48.02 61.43 54.36
w/o QDC 46.54 59.28 52.80
6 CFS 48.73 63.38 56.56
w/o QDC 48.00 63.10 56.29
CIS 47.34 61.30 54.55
w/o QDC 46.30 60.00 53.95
Time for One Response    Compared to previous RAG frameworks, DR-RAG also achieves better time optimization during the whole process. Other RAG frameworks may call LLMs multiple times before obtaining the answer, resulting in high computational cost. In fact, the inference time of LLMs is also a worthwhile optimization in the applications. It takes a lot of time to call LLMs once, and calling them multiple times presents a catastrophic challenge in terms of time overhead. Therefore, we attempt to design a small model with relatively few parameters to achieve better optimization rather than calling LLMs multiple times. In Table 2, compared to Adaptive-RAG, we have achieved an average 74.2% reduction in time overheads. Therefore, we can conclude that we can achieve better experimental efficiency and the time overhead makes DR-RAG valuable for practical engineering applications.

### v3

@@ Abstract
11footnotetext: Contribute equally during internshiping at Li Auto.
Retrieval-Augmented Generation (RAG) has recently demonstrated the performance of Large Language Models (LLMs) in the knowledge-intensive tasks such as Question-Answering (QA). RAG expands the query context by incorporating external knowledge bases to enhance the response accuracy. However, it would be inefficient to access LLMs multiple times for each query and unreliable to retrieve all the relevant documents by a single query. We have found that even though there is low relevance between some critical documents and query, it is possible to retrieve the remaining documents by combining parts of the documents with the query. To mine the relevance, a two-stage retrieval framework called Dynamic-Relevant Retrieval-Augmented Generation (DR-RAG) is proposed to improve document retrieval recall and the accuracy of answers while maintaining efficiency. Additionally, a compact classifier is applied to two different selection strategies to determine the contribution of the retrieved documents to answering the query and retrieve the relatively relevant documents. Meanwhile, DR-RAG call the LLMs only once, which significantly improves the efficiency of the experiment. The experimental results on multi-hop QA datasets show that DR-RAG can significantly improve the accuracy of the answers and achieve new progress in QA systems.
DR-RAG: Applying Dynamic Document Relevance to Retrieval-
Augmented Generation for Question-Answering
Zijian Hei∗,1,2, Weiling Liu∗,1,3, Wenjie Ou∗,1,4, Juyi Qiao1, Junming Jiao1, Guowen Song,122footnotemark: 2, Ting Tian,233footnotemark: 3, Yi Lin,433footnotemark: 3 1 Li Auto Inc.,2 Sun Yat-sen University,3 Northeastern University, China,4 Sichuan University songguowen@lixiang.com tiant55@mail.sysu.edu.cn, yilin@scu.edu.cn
@@ 1 Introduction
Large language models (LLMs) have recently made significant improvement in the field of Natural Language Processing (NLP), especially in text generation tasks Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023b); Anil et al. (2023); Ouyang et al. (2022); Touvron et al. (2023a). Although LLMs excel in various application scenarios, challenges remain regarding the accuracy and timeliness of the generated text, especially in real-time domains. LLMs with intrinsic parameter memories may generate inaccurate or even incorrect text when faced with up-to-date query Min et al. (2023); Mallen et al. (2022); Muhlgay et al. (2023). This issue, known as hallucination, occurs when the text generated by LLMs fails to align with real-world knowledge Ji et al. (2023); Zhang et al. (2023); Kwiatkowski et al. (2019). Therefore, Retrieval-Augmented Generation (RAG) frameworks have been proposed to improve the accuracy of generated text by combining relevant information from external knowledge base with query Arora et al. (2023); Lewis et al. (2020); Borgeaud et al. (2022). RAG has effectively demonstrated its superiority in knowledge-intensive tasks such as open-domain Question-Answering (QA) and has achieved new progress in the LLMs’ performance.
@@ 2 Method
Figure 2: An overview of DR-RAG. In step 1, we retrieve static-relevant documents (SR-Documents) due to high relevance with the query. Then we concatenate SR-Documents with the query to retrieve multiple dynamic-relevant documents (DR-Documents) in step 2. Finally, we select each of DR-Documents in turn to concatenate with the query and SR-Documents and feed them into the classifier to select the most relevant DR-Document.
@@ 2.2 DR-RAG
In this section, we will give a comprehensive description about the DR-RAG framework, a new two-stage retrieval method compared to traditional reranking methods NetEase Youdao (2023); Chen et al. (2024). From Fig. 2, we retrieve $k_{1}$ documents through SM method (first-retrieval stage) and employ a classifier $C$ to model the dynamic relevance between documents (selection process) to enhance the recall rate of the remaining $k_{2}$ documents. The classifier $C$ lies in assessing the dynamic relevance between documents to determine whether the information from the documents is crucial to answer the present query.
@@ 2.2.1 Query Documents Concatenation
As mentioned before, due to the low relevancy between dynamic-relevant documents and the query, the documents are difficult to be retrieved. Moreover, the only relevant information ‘spouse/wife’ between them is also obscured by the mixed information in the knowledge base because too many documents in $D$ will contain ‘spouse’. Therefore, Query Documents Concatenation (QDC) method aims to employ the sentence to match for more useful and relevant information. After the first-retrieval stage, we will obtain $k_{1}$ static-relevant documents and concatenate $\boldsymbol{q}$ with each document to form multiple <$\boldsymbol{q},\boldsymbol{d_{i}},i\in k_{1}$> pairs. Moreover, dynamic-relevant documents from $D$ can be retrieved by corresponding <$\boldsymbol{q},\boldsymbol{d_{i}},i\in k_{1}$> pair in the second-retrieval stage. As the case in Fig 2, when $\boldsymbol{q}$ and $\boldsymbol{d_{stat}^{*}}$ are concatenated, the query contains both the ‘Johan Ludvig Heiberg’ and the relationship ‘spouse/wife’, which is essentially similar to $\boldsymbol{d_{dyn}^{*}}$. Therefore, $\boldsymbol{d_{dyn}^{*}}$ is more clearly related to the query and thus easily retrieved. The whole process is:
$\begin{split}Cnt&=\{\}\\ \{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}&=\texttt{Retriever}(\boldsymbol{q})\\ Cnt&=Cnt\cup\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}\\ \boldsymbol{q}_{i}^{*}&=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d}_{i})\\ \{\boldsymbol{d}^{\prime}_{i,1},\ldots,\boldsymbol{d}^{\prime}_{i,k_{2}}\}&=\texttt{Retriever}(\boldsymbol{q}_{i}^{*})\\ Cnt&=Cnt\cup\{\boldsymbol{d}^{\prime}_{i,j}\mid\boldsymbol{d}^{\prime}_{i,j}\not\in Cnt\land first\}\\ answer&=\texttt{LLM}(\texttt{Concat}(\boldsymbol{q},Cnt))\end{split}$ (1)
where $k_{1}+k_{2}$ is equal to $k$. Retriever is a common SM method. $\boldsymbol{d}$ and $\boldsymbol{d^{\prime}}$ are the relevant document retrieved from $D$ in the first and second-retrieval stage. $Cnt$ is a context containing multiple documents. $Cnt=Cnt\cup\{\boldsymbol{d}^{\prime}_{i,j}\mid\boldsymbol{d}^{\prime}_{i,j}\not\in Cnt\land first\}$ means that for for a given $\boldsymbol{d^{\prime}}$, the first $\boldsymbol{d}^{\prime}_{i,j}$ in the second-retrieval stage that is not already part of $Cnt$ will be placed into $Cnt$. LLM is a large language model to obtain the answer. $answer$ is the output to answer the query.
@@ 2.2.2 Classifier for Selection
$\begin{split}\text{C}(\boldsymbol{q},\boldsymbol{d}^{*},\boldsymbol{d}^{*})&=positive\\ \text{C}(\boldsymbol{q},\boldsymbol{d}^{*},\boldsymbol{d}^{\Delta})&=negative\\ \text{C}(\boldsymbol{q},\boldsymbol{d}^{\Delta},\boldsymbol{d}^{\Delta})&=negative\\ \end{split}$ (2)
where C represents the classifier. $positive$ and $negative$ indicate whether the two documents are critical for the query.
Table 2: Results on different datasets with Llama3-8B as LLM. Adaptive Retrieval and Self-RAG conduct the retrieval module only under specific conditions (unpopular query entities or special retrieval tokens), so their time overhead is much less than other methods. We emphasize our results in bold.
@@
Classifier Inverse Selection    In this approach, we selectively exclude some irrelevant documents from the retrieved $k$ documents to minimize document redundancy. Specifically, after obtaining $k$ documents in stages, we pair them as <$\boldsymbol{q},\boldsymbol{d_{m}},\boldsymbol{d_{n}}$> and get $C_{k}^{2}$ pairs. The pairs, with the current query $\boldsymbol{q}$, are collectively fed into the classifier $C$. Similarly, when the classification result of a document and the remaining $k$-1 documents is negative, then we consider the document as redundant and should be removed. The whole process is:
$\begin{split}Cnt&=Cnt\cup\{\boldsymbol{d}^{\prime}_{i,j}\mid\boldsymbol{d}^{\prime}_{i,j}\not\in Cnt\land\text{$first$}\}\\ P_{i,j}&=\begin{cases}1&\text{if }\exists i,C(\boldsymbol{q},\boldsymbol{d^{\prime}}_{i,j},\boldsymbol{d}_{i})=positive\\ 0&\text{$otherwise$}\end{cases}\\ Cnt&=Cnt-\{\boldsymbol{d^{\prime}}_{i,j}\mid P_{i,j}=0\}\\ \text{$answer$}&=\texttt{LLM}(\texttt{Concat}(\boldsymbol{q},Cnt))\end{split}$ (3)
where - represents complement. $Cnt=Cnt-\{\boldsymbol{d^{\prime}}_{i,j}\mid P_{i,j}=0\}$ means $\boldsymbol{d^{\prime}}_{i,j}$ in the second-retrieval stage is classified as $negative$ combined with all $\boldsymbol{di}$ in the first-retrieval stage, then $\boldsymbol{d^{\prime}}_{i,j}$ will be removed.
Classifier Forward Selection    Unlike the CIS method, CFS method aims to remove the irrelevant dynamic-relevant documents in the second-retrieval stage. To achieve this goal, we search for a document $\boldsymbol{d_{n}}$ from $D$ according to the <$\boldsymbol{q},\boldsymbol{d_{m}}$> pair, and feed both the query and documents into $C$. When the classification result is negative, we will exclude the dynamic-relevant document in the current retrieved documents, and search for the next dynamic-relevant document which can be classified as positive with m. The whole process is:
$\begin{split}P_{i,j}&=\begin{cases}1&\text{if }C(\boldsymbol{q},\boldsymbol{d}_{i},\boldsymbol{d}^{\prime}_{i,j})=positive\\ 0&\text{$otherwise$}\end{cases}\\ Cnt&=Cnt\cup\{\boldsymbol{d}^{\prime}_{i,j}\mid P_{i,j}=1\land first\}\\ \text{$answer$}&=\texttt{LLM}(\texttt{Concat}(\boldsymbol{q},Cnt))\end{split}$ (4)
where $Cnt=Cnt\cup\{\boldsymbol{d}^{\prime}_{i,j}\mid P_{i,j}=1\land first\}$ means that for a given $\boldsymbol{d_{i}}$, the first $\boldsymbol{d}^{\prime}_{i,j}$ in the second-retrieval stage classified as $positive$ combined with $\boldsymbol{d_{i}}$ will be considered as dynamic-relevant document and placed into $Cnt$.
@@ 3 Experiment Settings
The experimental details will be described in this section. Due to space constraints, the descriptions of implementation details, retrieval strategy and baseline can seen in Appendix A.1, A.2 and A.3.
@@ 3.1 Dataset
Table 3: Results on different LLMs and strategies compared to Adaptive-RAG. We set gpt-3.5-turbo and Llama3-8b as the base LLM. We emphasize our best results in bold. Top-k means the total number of retrieved documents.
@@
Llama3-8B Query Document Concatenation 25.10 37.10 32.00 46.05 59.98 53.09 45.79 54.62 49.70
@@
Table 4: Ablation study on HotpotQA by Llama3-8B.
top-k LLMs EM F1 Acc
3 CFS 48.71 62.42 55.26
w/o QDC 46.26 59.37 52.60
CIS 44.88 57.05 50.94
w/o QDC 44.82 57.00 50.88
4 CFS 48.30 62.81 55.22
w/o QDC 47.78 61.63 54.73
CIS 48.02 61.43 54.36
w/o QDC 46.54 59.28 52.80
6 CFS 48.73 63.38 56.56
w/o QDC 48.00 63.10 56.29
CIS 47.34 61.30 54.55
w/o QDC 46.30 60.00 53.95
@@ 4 Results and Analysis
Table 5: Recall rate and actual numbers under different retrieval strategies. Actual numbers represents the actual numbers of documents that we feed into LLMs. A smaller number means fewer redundant documents.
@@ 4.1 Main Results
Table 2 and 3 present the performance of DR-RAG in answering multi-hop query, and highlight the advantages of our approach compared to the sota RAG framework Jeong et al. (2024); Asai et al. (2024) across multiple metrics, which is in line with our expectations. Table 5 shows the performance of DR-RAG across various retrieval strategies.
As shown in Table 2, when retrieving the same $k$ documents, DR-RAG can achieve a higher recall rate and a higher percentage of correct answers. From the results, DR-RAG achieves better performance than other baseline RAG frameworks (self-RAG and Adaptive-RAG) on all three metrics. Moreover, DR-RAG is also less than other RAG frameworks in terms of the number of LLMs responses and the time consumed in QA systems.
@@ 4.2 Analysis
Ablation Study    We propose a two-stage retrieval and classifier selection strategies to mine the dynamic relevance of documents. As shown in Table 3, we apple two classification methods based on QDC, and the experimental results have achieved further improvement. Table 4 shows the comparison of the effect of DR-RAG with and without QDC. Quantitatively, CIS and CFS can improve DR-RAG’s performance by 2.3% and 4.7% on Acc metric against QDC, while DR-RAG reduces performance by 1.1% and 0.7% on Acc metric without QDC. The results demonstrate that the two strategies are able to efficiently extract document relevance and achieve more accurate answers.
Table 6: Results of different classifier on HotpotQA dataset as Llama3-8B.
top-$k$ Classifier EM F1 Acc
3 Bigbird-base(125M) 48.71 62.42 55.26
Bigbird-large(355M) 49.20 62.89 55.55
Longformer(147M) 49.00 62.80 55.34
4 Bigbird-base(125M) 48.30 62.81 55.22
Bigbird-large(355M) 49.12 63.70 56.00
Longformer(147M) 50.27 64.41 56.90
6 Bigbird-base(125M) 48.73 63.38 56.56
Bigbird-large(355M) 49.02 63.58 56.52
Longformer(147M) 50.08 64.69 57.30
Table 7: Results of 500 samples sampled on HotpotQA dataset based on gpt-3.5-turbo and gpt-4-turbo.
top-$k$ LLMs EM F1 Acc
3 gpt-3.5-turbo 43.80 58.83 55.00
gpt-4-turbo 51.60 68.93 67.20
4 gpt-3.5-turbo 45.00 60.55 57.40
gpt-4-turbo 53.40 70.75 69.80
6 gpt-3.5-turbo 48.20 63.83 63.60
gpt-4-turbo 59.20 75.43 77.20
Effects of Classifier and LLM    Compared to gpt-3.5-turbo, gpt-4-turbo with better document comprehension has the ability to accurately capture the critical information to answer a query. As for textual responses, gpt-4-turbo generates responses of higher quality and more accurate content. Quantitatively, as shown in Table 7, gpt-4-turbo improve by an average of 9.07%, 10.63%, and 12.73% against gpt-3.5-turbo on three metrics. As shown in Table 6, when switching to different kinds or sizes of classifiers, the difference in the metrics is negligible (the extreme difference of EM, F1, and Acc is less than 2%), which suggests that our approach is applicable to different classifiers and that the classifier has little impact on our framework.
Effects of Recall Rate    The ability of LLMs to answer domain-specific query correctly almost depends on whether all the necessary information is included in the prompt context. When relevant information is missing, it is difficult for LLMs with the hallucination problem to accurately answer the query. Table 8 illustrates the answers of the query with and without sufficient information provided to LLMs. As seen in Table 5, in 2Wiki, our retrieval strategy already achieves a recall rate of 98% when top-$k$ is 6. When we feed enough relevant information into LLMs, the accuracy of their answers can be improved accordingly. CFS method achieves higher recall rate by 26.4% and 8.6% than BM25 and SM methods, respectively, which proves the feasibility of DR-RAG.
@@
Increase Recall with Lower Documents    In CFS method, it seems impossible to find a match for every <$\boldsymbol{q},\boldsymbol{d}$> pair in the second-retrieval stage because the documents we need have been retrieved. Therefore, there will be cases where the total number of our retrieved documents is less than $k$. For instance, in the HotpotQA dataset, when $k$ is set to 6, the average number of documents actually provided to LLMs is 5.35, thereby reducing irrelevant information to some extent. CFS method in Table 5 achieves a higher recall rate while retrieving fewer actual numbers of documents compared to QDC method. CFS method yields higher scores across the three metrics in our experiments and achieves more significant retrieval capabilities with lower redundant inputs than other methods.
Figure 3: QA performance (F1) and time for different RAG frameworks. We use the GPT-3.5-turbo as the base LLM on the multi-hop QA datasets (MuSiQue, HotpotQA and 2Wiki).
Time for One Response    Compared to previous RAG frameworks, DR-RAG also achieves better time optimization during the whole process. Other RAG frameworks may call LLMs multiple times, resulting in high computational cost. In fact, the inference time of LLMs is also a worthwhile optimization in the applications. It takes a lot of time to call LLMs once, and calling them multiple times presents a catastrophic challenge in terms of time overhead. Therefore, we attempt to design a small model with relatively few parameters to achieve better optimization rather than calling LLMs multiple times. In Fig. 3 and Table 2, compared to Adaptive-RAG, we have achieved an average 74.2% reduction in time overheads. Therefore, we can conclude that we can achieve better experimental efficiency and the time overhead makes DR-RAG valuable in applications.